Status Report

NASA Space Station On-Orbit Status 19 Dec 2003

By SpaceRef Editor
December 19, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 19 Dec 2003

All ISS systems continue to function nominally, except as noted previously or below.  The crew’s 62nd day in space (60 aboard ISS).

Early in the morning, before breakfast, FE Kaleri concluded Part 3 of his current MBI-8 “Profilaktika” (countermeasures) fitness session.  Today he conducted the blood tests, which determine lactate and creatine kinase levels in the blood with the AccuSport equipment, followed by the physical exercise session on the TVIS treadmill.  CDR Foale assisted as CMO (crew medical officer) as required.  [The TVIS test is identical to the MO-3 test performed on the treadmill in idling mode with free choice of speeds within a certain specified range.  In addition to the nominal test procedure, MBI-8/Part 3 calls for the use of the TEEM-100M gas analyzer during the test, the blood lactate measurements, and subjective evaluation of physical exertion levels during the test.  At the end of the creatine kinase tests, the results were logged, copied from Cardiocassette-2000 recording and reported to the ground.]

Both crewmembers underwent their second 30-min. CBT (computer-based training with video and audio) exercise to refresh their Crew Medical Officer (CMO) proficiency in emergency medical operations.  [Objective of today’s CBT, scheduled regularly once a month, was to review information regarding nasal airway, equipment & operations with the suction device, the ILMA (intubating laryngeal mask airway), and performing a cricothyrotomy (making an incision under the Adam’s apple to bypass an airway obstruction).  A debriefing with the Surgeon was scheduled afterwards.]

Alex Kaleri worked in the FGB, to replace a standardized switching device (UKP) in the onboard complex control system (SUBK). [Seven of these relay switches are used by the SUBK to control onboard systems and components in the FGB, in one of three modes: (a) automatic (using ground commands via the Komparus system, commands from the Russian and U.S. segments, and commands generated by the SM computer), (b) program relay (backup, using Komparus and SUBK relay components), and (c) command relay (commands from the ground via the Komparus comm portal or transmitted from the SM.)]

Kaleri took digital photographs of the RAMs (radiation area monitors), deployed throughout the ISS.  The images are to be downlinked at next earliest convenience. [There are 16 units (SM/6, Node/3, A/L/2, Lab/5), and Sasha took two photos of each, one a close-up, the other showing the RAM in relation to other panels.]

Both crewmembers in turn took the periodic on-orbit hearing assessment (O-OHA) test, a NASA EHS (environmental health systems) examination to assess the efficacy of acoustic countermeasures.  O-OHA, deferred from 12/15, was last done on 11/11. [The O-OHA test involves minimum audibility measurements for each ear over a wide range of frequencies and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter).  To conduct the testing, the experimenter is supported by special “EarQ” software on the MEC (medical equipment computer).  The baseline test is required for about Flight Day 14 for each new Expedition and is then performed once per month.]

After the ground had powered up the HRF (Human Research Facility), and with its laptop working again, Mike Foale transferred the experiment results of the FOOT (Foot/Ground Reaction Forces during Space Flight) run conducted on 12/3 from PCMCIA memory card to the PC for later downlink.  HRF was then deactivated again.

The CDR completed the regular computer task of transferring data files from the physical exercise equipment to the MEC via memory card and RED log entries, for downlink on OCA comm.  Afterwards, he also performed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, afterwards deleting them on the HRM.   

Mike performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).

Today’s planned 1-hr. repress of the cabin atmosphere with fresh O2 from Progress-248 (12P) storage was not required and deleted.  [Since MCC-H regularly wants to have the MCA (Major Constituents Analyzer) active during the represses for partial pressure readings, Houston requires a 14-hr. advance notice of the repress event planned by Moscow.  This has not been observed or necessary in the past.]

Foale performed the regular status checkup of Lab payloads, while the FE conducted the daily routine maintenance of the SOZh life support system and prepared the prepared the IMS inventory “delta” file.  Alex also conducted the regular 20-min. tag-up with ground specialists via S-band to discuss the weekly IMS (inventory management system) status update.

At ~10:10am EST, the CDR broke out Christmas decorations from stowage (Santa hats, presents from families) and prepared cabin interior and themselves for today’s PAO event, a special message with Christmas greetings downlinked at 11:08am for use on NASA TV.  Afterwards, the crew sent down a second special greeting, this time intended for New Year.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for some viewing angles, were Lower Yangtze smog, China (Dynamic event: Smog event in the lower Yangtze and Yellow river valleys is making the news again.  Handheld images were requested to illustrate a different perspective on these increasingly large haze masses), Berlin, Germany (pass over the city.  Looking a touch left for the city center), Aral Sea, Kazakhstan (good pass across the SW corner of the sea.  General views requested to complement past detailed ISS/CEO images), Ganges valley aerosols (the next smog accumulation event appears to be starting, with visibilities down to 2 miles at cities in the Ganges valley.  Looking half left for oblique views down the valley.  The last smog event made world news.  Scientists say that handheld images are a prime method of helping to visualize these masses), Subglacial rivers, SW Libya (400mm-lens: Mapping swath right of track for ~30 secs was requested to document sub-glacial river beds.  ISS/CEO images are aimed at reconstructing the broader, regional patterns of these sinuous, ancient river beds), Lagos, Nigeria (looking a touch right on either side of the coastal lagoon), Miami, FL (the Everglades wetland is creating an ever stronger barrier to westward urban growth of Miami.  Patterns of city growth reflect this barrier), and Caracas, Venezuela (Caracas is finally starting to appear in view as the equatorial cloud masses shift further south.  Looking a touch left, on the coast, for this major port through which a major proportion of US oil is exported).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.8; ppO2 (mmHg) — 142.7; ppCO2 (mmHg)  4.6
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 19.8.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C)  21.7.
  • Node:  Pressure (mmHg) — 743.80; temperature (deg C) — 24.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 746.9; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 747.04; temperature (deg C) — 24.5; shell heater temp (deg C) — 24.4, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.5
  • PMA-2:  Shell heater temp (deg C) — 11.7

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking, “night glider”/”sun slicer” drag reduction mode). 
  • SM batteries:  Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting, off-line; all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 3680 kg (8113 lb) as of 12/11/03  [SM(755) + FGB(2573) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -89.9 deg, pitch: -8.9 deg, roll: 1.8 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1/LEE B, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:50am EST [= epoch]):

  • Mean altitude — 371.2 km
  • Apogee — 375.2 km
  • Perigee — 367.2 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0005881
  • Orbits per 24-hr. day — 15.66
  • Solar Beta angle —  -71.6, decreasing
  • Mean altitude loss in last 24 hours — 70
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28998
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.